OpenStudy (anonymous):

Solve. Show work please for a more clear understanding 2^((x+1))=8^((x-3))

OpenStudy (asnaseer):

notice that \(8=2^3\) - make use of that fact

OpenStudy (anonymous):

What do you mean?

OpenStudy (asnaseer):

rewrite:\[8^{x-3}\]by using the fact that:\[8=2^3\]

OpenStudy (asnaseer):

@ksaimouli - the aim here is to teach - not just give answers out

OpenStudy (ksaimouli):

@rmwdrika who is better

OpenStudy (anonymous):

I just need that one example. I didn't ask for my whole assignment

OpenStudy (anonymous):

@ksaimouli can you please post that again?

OpenStudy (asnaseer):

@rmwdrika - this site is not intended for just getting answers. We encourage members to learn the subject matter.

OpenStudy (anonymous):

I asked for the work as well. I just needed that one example

OpenStudy (asnaseer):

yes - and I am trying to guide you to the answer.

OpenStudy (asnaseer):

so first try and rewrite \(8^{x-3}\) by replacing the 8 with \(2^3\)